OpenStudy (anonymous):

Describe and provide an example of a system of equations with one solution, No solution, and Infinite solutions.

OpenStudy (anonymous):

x=y^2 x=-1 Has no solution x=y^2 x=-y^2 Has one solution (0,0) x=y^2 sqrt x=y Has infinite solutions